# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v1

@@ Appendix B Rephrased Prompts
System
You’re participating in a game involving $N$ participants and it spans across $K$ rounds.
@@
User
The outcomes of the game for Round $I$ are as follows:
@@
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
@@
User
Commencing with round $I$.
@@
System
You’re engaged in a game, involving $N$ participants across $K$ rounds.
@@
User
Results of Round $I$ Game:
@@
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
@@
User
The commencement of round $I$ is now.
@@
System
You’re involved in a game which brings $N$ participants together for $K$ rounds.
@@
User
The outcomes for Round $I$ are as follows:
@@
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
@@
User
The commencement of round $I$ is now.
@@
System
You will be engaging in a game that is played over $K$ rounds and includes a total of $N$ players.
@@
User
Results of the Game for Round $I$:
@@
Assistant
{“chosen_number”: “$C_{IJ}$”}
User
[Congratulation you won]/[Unfortunately you lost].
@@
User
The commencement of round $I$ is now.
@@ Appendix C Rescale Method for Raw Scores
$\begin{split}S_{1}&=\begin{cases}\frac{MAX-S_{1}}{MAX-MIN}*100,&R<1\\ \frac{\lvert 2S_{1}-(MAX-MIN)\rvert}{MAX-MIN}*100,&R=1\\ \frac{S_{1}}{MAX-MIN}*100,&R>1\end{cases},\\ S_{2}&=\frac{\max(R,1-R)-S_{2}}{\max(R,1-R)}*100,\\ S_{3}&=\frac{G-S_{3}}{G}*100,\\ S_{4}&=\begin{cases}\frac{T-S_{4}}{T}*100,&R\leq 1\\ \frac{S_{4}}{T}*100,&R>1\end{cases},\\ S_{5}&=S_{5}*100,\\ S_{6}&=100-S_{6},\\ S_{7}&=S_{7}*100,\\ S_{8}&=\frac{2*G-S_{8P}}{2*G}*50+S_{8V}*50.\\ \end{split}$ (1)
@@ Appendix D More Quantitative Results
Public Goods Game $58.8$ $74.7$ $54.3$ $62.1$ $56.1$ $61.2_{\pm 8.1}$
Diner’s Dilemma $96.0$ $96.5$ $100.0$ $93.5$ $100.0$ $97.2_{\pm 2.8}$
Sealed-Bid Auction $88.3$ $87.0$ $86.0$ $87.9$ $84.2$ $86.7_{\pm 1.6}$
@@
Public Goods Game $84.8$ $89.3$ $82.2$ $82.0$ $63.6$ $58.8$ $76.7_{\pm 12.5}$
Diner’s Dilemma $100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $96.0$ $99.3_{\pm 1.6}$
Sealed-Bid Auction $88.1$ $86.7$ $87.9$ $89.6$ $90.4$ $88.3$ $88.5_{\pm 1.3}$
@@
Public Goods Game $58.8$ $73.4$ $54.9$ $49.8$ $75.8$ $62.5_{\pm 11.5}$
Diner’s Dilemma $96.0$ $96.5$ $100.0$ $43.0$ $81.5$ $83.4_{\pm 23.7}$
Sealed-Bid Auction $88.3$ $89.6$ $89.1$ $89.7$ $80.5$ $87.4_{\pm 3.9}$
@@
Public Goods Game $58.8$ $43.9$ $67.6$ $62.6$ $74.4$
Diner’s Dilemma $69.0$ $17.5$ $100.0$ $82.5$ $53.0$
Sealed-Bid Auction $88.3$ $95.4$ $88.5$ $90.0$ $87.6$

### v2

@@ Appendix B Rephrased Prompts
System You’re participating in a game involving $N$ participants and it spans across $K$ rounds.
@@
User The outcomes of the game for Round $I$ are as follows:
@@
Assistant {“chosen_number”: “$C_{IJ}$”}
User [Congratulation you won]/[Unfortunately you lost].
@@
User Commencing with round $I$.
@@
System You’re engaged in a game, involving $N$ participants across $K$ rounds.
@@
User Results of Round $I$ Game:
@@
Assistant {“chosen_number”: “$C_{IJ}$”}
User [Congratulation you won]/[Unfortunately you lost].
@@
User The commencement of round $I$ is now.
@@
System You’re involved in a game which brings $N$ participants together for $K$ rounds.
@@
User The outcomes for Round $I$ are as follows:
@@
Assistant {“chosen_number”: “$C_{IJ}$”}
User [Congratulation you won]/[Unfortunately you lost].
@@
User The commencement of round $I$ is now.
@@
System You will be engaging in a game that is played over $K$ rounds and includes a total of $N$ players.
@@
User Results of the Game for Round $I$:
@@
Assistant {“chosen_number”: “$C_{IJ}$”}
User [Congratulation you won]/[Unfortunately you lost].
@@
User The commencement of round $I$ is now.
@@ Appendix C Rescale Method for Raw Scores
$\begin{split}S_{1}&=\begin{cases}\frac{MAX-S_{1}}{MAX-MIN}*100,&R<1\\ \frac{\lvert 2S_{1}-(MAX-MIN)\rvert}{MAX-MIN}*100,&R=1\\ \frac{S_{1}}{MAX-MIN}*100,&R>1\end{cases},\\ S_{2}&=\frac{\max(R,1-R)-S_{2}}{\max(R,1-R)}*100,\\ S_{3}&=\frac{G-S_{3}}{G}*100,\\ S_{4}&=\frac{T-S_{4}}{T}*100,\\ S_{5}&=(1-S_{5})*100,\\ S_{6}&=\frac{S_{6}}{\max_{ij}v_{ij}}*100,\\ S_{7}&=S_{7}*100,\\ S_{8}&=\frac{2*G-S_{8P}}{2*G}*50+S_{8V}*50.\\ \end{split}$ (1)
@@ Appendix D More Quantitative Results
Public Goods Game $41.3$ $25.4$ $45.7$ $38$ $44.0$ $38.8_{\pm 8.1}$
Diner’s Dilemma $4.0$ $3.5$ $0.0$ $6.5$ $0.0$ $2.8_{\pm 2.8}$
Sealed-Bid Auction $6.5$ $6.5$ $4.6$ $5.5$ $4.4$ $5.5_{\pm 1.0}$
@@
Public Goods Game $15.3$ $10.8$ $17.9$ $18.0$ $36.5$ $41.3$ $23.3_{\pm 12.5}$
Diner’s Dilemma $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $4.0$ $0.7_{\pm 1.6}$
Sealed-Bid Auction $5.6$ $6.0$ $4.6$ $4.5$ $5.8$ $6.5$ $5.5_{\pm 0.8}$
@@
Public Goods Game $41.3$ $26.6$ $45.2$ $50.2$ $24.2$ $37.5_{\pm 11.5}$
Diner’s Dilemma $4.0$ $3.5$ $0.0$ $57.0$ $18.5$ $16.6_{\pm 23.7}$
Sealed-Bid Auction $6.5$ $4.6$ $5.7$ $2.6$ $5.9$ $5.0_{\pm 1.6}$
@@
Public Goods Game $41.3$ $56.1$ $32.4$ $37.4$ $25.6$
Diner’s Dilemma $31.0$ $82.5$ $0.0$ $17.5$ $47.0$
Sealed-Bid Auction $6.5$ $1.4$ $7.4$ $4.0$ $5.8$